OpenStudy (anonymous):

In a 45°- 45°- 90° right triangle, the length of the hypotenuse is 15. How long are the legs?

OpenStudy (anonymous):

use trig to find the sides. sin=opp/hyp, cos=adj/hyp, and tan=opp/adj. You are given your hypotenuse (hyp) of 15 and you know your angles, sin(45) cos(45), and solve for your missing pieces.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Do you need an example or does that make sense?

OpenStudy (anonymous):

mabe an example will help a little

OpenStudy (anonymous):